# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 6 Conclusion
S2.1 Expressions for $\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$
In the reparameterisation gradient given in Equation (S9), $\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$. We write that $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+(\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s})\text{vech}(\textbf{B})+\textbf{d}\circ\boldsymbol{\epsilon}$, where $\otimes$ represents the Kronecker product, and $\textbf{I}_{s}$ is the identity matrix of size $s$. It can be shown that $\nabla_{\boldsymbol{\xi}}\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})=-(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\boldsymbol{\xi}-\boldsymbol{\mu})$,
$\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\mu}}=\textbf{I}_{s}\\ ~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\text{vech}(\textbf{B})}=\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s}.$ (S10)
The derivatives of the lower bound with respect to variational parameters are:
$\begin{split}\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}~{}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})],\end{split}$ (S11)
$\begin{split}\nabla_{\text{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}],\end{split}$ (S12)
@@
$\begin{split}\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\text{diag}(\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top})],\end{split}$ (S13)
@@
S2.2 Expressions for $\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})=\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\xi})$
The term $h(\boldsymbol{\xi})$ represents the likelihood multiplied by the prior distribution (see Section S2). For different SEMs, the expressions for $h(\boldsymbol{\xi})$ are given in Table S2. Table S3 shows the prior distributions used for the model parameters in this study. In this section, we compute the derivatives of $\log h(\boldsymbol{\xi})$ with respect to the model parameter vector $\boldsymbol{\xi}$, denoted as $\nabla_{\boldsymbol{\xi}}\log h(\boldsymbol{\xi})$. These derivatives are necessary for computing the reparameterisation gradient, as outlined in Equation (S9), for each SEM.
@@
Table S3: Prior distributions of model parameters
S2.2.1 Derivatives for SEM-Gau
For SEM-Gau, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
@@
$\begin{split}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)&=-(\textbf{W}^{\top}+\textbf{W})+2\rho\textbf{W}^{\top}\textbf{W},\\ \left(\frac{\partial\rho}{\partial\rho^{\prime}}\right)&=\frac{2e^{\rho^{\prime}}}{(1+e^{\rho^{\prime}})^{2}}.\\ \end{split}$
@@
For SEM-t, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\nu$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
@@
$\begin{split}\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\nu^{\prime}}&=n\left(\frac{\partial~{}\textrm{log}~{}\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}\right)-\frac{e^{\nu^{\prime}}}{2}\sum_{i=1}^{n}\left(\tau_{i}^{\prime}+e^{-\tau_{i}^{\prime}}\right)-\frac{ne^{\nu^{\prime}}}{2}\textrm{log}\left(\frac{2}{e^{\nu^{\prime}}+3}\right)+\frac{ne^{\nu^{\prime}}}{2}\\ &-\frac{\nu^{\prime}}{\sigma^{2}_{\nu^{\prime}}},\\ \end{split}$ (S22)
@@
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $i^{th}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
@@
where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{th}$ element of the vector Ar.
@@
For YJ-SEM-Gau, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
@@
where $\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}=\left(\frac{\partial t_{\gamma}(y_{1})}{\partial\gamma},\frac{\partial t_{\gamma}({y}_{2})}{\partial\gamma},\ldots,\frac{\partial t_{\gamma}({y}_{n})}{\partial\gamma}\right)^{\top}$, with the $i^{th}$ elements of the vector is equal to
@@
For YJ-SEM-t, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
@@
$\begin{split}\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\nu^{\prime}}&=n\left(\frac{\partial~{}\textrm{log}~{}\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}\right)-\frac{e^{\nu^{\prime}}}{2}\sum_{i=1}^{n}\left(\tau_{i}^{\prime}+e^{-\tau_{i}^{\prime}}\right)-\frac{ne^{\nu^{\prime}}}{2}\textrm{log}\left(\frac{2}{e^{\nu^{\prime}}+3}\right)+\frac{ne^{\nu^{\prime}}}{2}\\ &-\frac{\nu^{\prime}}{\sigma^{2}_{\nu^{\prime}}},\\ \end{split}$ (S33)
@@
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $i^{th}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
@@
where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{th}$ element of the vector Ar.
@@
The adaptive learning rates (step sizes) for the VB algorithms are calculated using the ADADELTA algorithm (Zeiler,, 2012). The ADADELTA algorithm is now briefly described. Different step sizes are used for each element in variational parameters $\boldsymbol{\lambda}$.
The update for the $i^{th}$ element of $\boldsymbol{\lambda}$ is
$\boldsymbol{\lambda}^{(t+1)}_{i}=\boldsymbol{\lambda}^{(t)}_{i}+\Delta\boldsymbol{\lambda}^{(t)}_{i},$ (S36)
where, the step size $\Delta\boldsymbol{\lambda}^{(t)}_{i}$ is $a_{i}^{(t)}g_{\boldsymbol{\lambda}_{i}}^{(t)}$. The term $g_{\boldsymbol{\lambda}_{i}}^{(t)}$ denotes the $i^{th}$ component of $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$ and $a_{i}^{(t)}$ is defined as:
$\mathcal{a}_{i}^{(t)}=\sqrt{\frac{E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t-1)}+\alpha}{E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}+\alpha}},$ (S37)
where $\alpha$ is a small positive constant, $E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda_{i}}}^{2}\right)^{(t)}$ and $E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}$ are decayed moving average estimates of ${\boldsymbol{\Delta}{\boldsymbol{\lambda_{i}}}^{(t)}}^{2}$ and ${g_{\boldsymbol{\lambda}_{i}}^{(t)}}^{2}$, defined by
$E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t)}=\upsilon E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t-1)}+(1-\upsilon){\boldsymbol{\Delta}{\boldsymbol{\lambda_{i}}}^{(t)}}^{2},$ (S38)
@@
$E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}=\upsilon E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t-1)}+(1-\upsilon){g_{\boldsymbol{\lambda}_{i}}^{(t)}}^{2},$ (S39)
@@
where the variable $\upsilon$ is a decay constant. We use the default tuning parameter choices $\alpha=10^{-6}$ and $\upsilon=0.95$, and initialize $E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(0)}=E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(0)}=0$.
S4 Additional Derivation of HVB algorithms for estimating SEMs with missing data
S4.1 Derivation of the reparameterisation gradient for HVB algorithm
@@
Since $\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda})$ as shown in Equation (16) of the main paper. Then, the reparameterisation gradient of $\mathcal{L}$ is the same as that of $\mathcal{L}^{0}$, given by

### v2

@@ 9 Conclusion
S2.1 Derivation of $\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$, $\nabla_{\boldsymbol{\xi}}\log q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ and $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$
In the reparameterisation gradient of the ELBO given in Equation (S9), $\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$. We write that $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+(\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s})\text{vech}(\textbf{B})+\textbf{d}\circ\boldsymbol{\epsilon}$}, where $\otimes$ represents the Kronecker product, and $\textbf{I}_{s}$ is the identity matrix of size $s$. It can be shown that $\nabla_{\boldsymbol{\xi}}\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})=-(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\boldsymbol{\xi}-\boldsymbol{\mu})$},
$\frac{\partial u(\boldsymbol{\zeta},\boldsymbol{\lambda})}{\partial\boldsymbol{\mu}}=\textbf{I}_{s}\\ ,~~~~~~~~~~\frac{\partial u(\boldsymbol{\zeta},\boldsymbol{\lambda})}{\partial\text{vech}(\textbf{B})}=\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s}\\ ,~~~\textrm{and}~~\frac{\partial u(\boldsymbol{\zeta},\boldsymbol{\lambda})}{\partial\textbf{d}}=\boldsymbol{\epsilon}\\ .$ (S10)
The derivatives of the ELBO with respect to variational parameters are:
$\begin{split}\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}~\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})],\end{split}$ (S11)
$\begin{split}\nabla_{\text{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}],\end{split}$ (S12)
@@
$\begin{split}\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\zeta}}}[\text{diag}(\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top})],\end{split}$ (S13)
@@
S2.2 Derivation of $\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})=\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\xi})$
The term $h(\boldsymbol{\xi})$ represents the likelihood multiplied by the prior distribution (see Section S2). For different SEMs, the expressions for $h(\boldsymbol{\xi})$ are given in Table S2. In this section, we present the derivatives (gradients) of $\log h(\boldsymbol{\xi})$ with respect to the model parameter vector $\boldsymbol{\xi}$, denoted as $\nabla_{\boldsymbol{\xi}}\log h(\boldsymbol{\xi})$. These gradients are required to construct unbiased estimates of the reparameterisation gradient.
To map the parameters $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$ onto the real line, we apply the following transformations: $\omega^{\prime}=\text{log}(\sigma^{2}_{\textbf{e}})$, $\rho^{\prime}=\text{log}(1+\rho)-\text{log}(1-\rho)$, $\nu^{\prime}=\text{log}(\nu-3)$, and $\gamma^{\prime}=\text{log}(\gamma)-\text{log}(2-\gamma)$. Prior distributions are then specified on the transformed parameter space. Table S3 shows the prior distributions used for the model parameters.
Table S3: Prior distributions of model parameters
@@
S2.2.1 Gradients for SEM-Gau
For SEM-Gau, we set $\boldsymbol{\xi}=\boldsymbol{\phi}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\mathbf{e}},\rho)^{\top}$, so that $h(\boldsymbol{\xi})=p(\mathbf{y}\mid\boldsymbol{\phi})\,p(\boldsymbol{\phi})$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
@@
[…]
@@
For the SEM-t, we set $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$, so that $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$}. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\nu$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
@@
[…]
@@
The derivative of $\text{log}~h(\boldsymbol{\xi})$ with respect to $i^{\textrm{th}}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
@@
where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{\textrm{th}}$ element of the vector Ar.
@@
For the YJ-SEM-Gau, we set $\boldsymbol{\xi}=\boldsymbol{\phi}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$, so that $h(\boldsymbol{\xi})=p(\mathbf{y}\mid\boldsymbol{\phi})\,p(\boldsymbol{\phi})$}. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\gamma$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
@@
where $\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}=\left(\frac{\partial t_{\gamma}(y_{1})}{\partial\gamma},\frac{\partial t_{\gamma}({y}_{2})}{\partial\gamma},\ldots,\frac{\partial t_{\gamma}({y}_{n})}{\partial\gamma}\right)^{\top}$, with the $i^{\textrm{th}}$ elements of this vector is equal to
@@
For the YJ-SEM-t, we set $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$, so that $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
@@
[…]
@@
The derivative of $\text{log}~h(\boldsymbol{\xi})$ with respect to $i^{\textrm{th}}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
@@
where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{\textrm{th}}$ element of the vector Ar.
@@
To ensure stable and efficient convergence of all VB algorithms presented in this paper, we use the ADADELTA algorithm (Zeiler,, 2012) to compute adaptive learning rates $\mathbcal{a}^{(t)}$. ADADELTA assigns different step sizes to each element of the variational parameter vector $\boldsymbol{\lambda}$. A brief description of the ADADELTA algorithm is provided below.
The update for the $i^{\textrm{th}}$ element of $\boldsymbol{\lambda}$, $\lambda_{i}$ is
${\lambda}^{(t+1)}_{i}={\lambda}^{(t)}_{i}+\Delta{\lambda}^{(t)}_{i}$, (S36)
where, the step size $\Delta{\lambda}^{(t)}_{i}$ is $a_{i}^{(t)}g_{{\lambda}_{i}}^{(t)}$. The term $g_{{\lambda}_{i}}^{(t)}$ denotes the $i^{th}$ component of $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$ and $a_{i}^{(t)}$ is defined as:
$\mathcal{a}_{i}^{(t)}=\sqrt{\frac{\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(t-1)}+\alpha}{\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t)}+\alpha}},$ (S37)
where $\alpha$ is a small positive constant, $\mathbb{E}\left({\Delta}_{{\lambda_{i}}}^{2}\right)^{(t)}$ and $\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t)}$ are decayed moving average estimates of ${{\Delta}{{\lambda_{i}}}^{(t)}}^{2}$ and ${g_{{\lambda}_{i}}^{(t)}}^{2}$, defined by
$\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(t)}=\upsilon\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(t-1)}+(1-\upsilon){{\Delta}{{\lambda_{i}}}^{(t)}}^{2},$ (S38)
@@
$\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t)}=\upsilon\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t-1)}+(1-\upsilon){g_{{\lambda}_{i}}^{(t)}}^{2},$ (S39)
@@
where the variable $\upsilon$ is a decay constant. We use the default tuning parameter choices $\alpha=10^{-6}$ and $\upsilon=0.95$, and initialise $\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(0)}=\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(0)}=0$.
S4 Selection model factorisation for SAR models
This section outlines how the selection model factorisation for SEMs, discussed in Section 3.2 of the main paper, can be straightforwardly extended to other simultaneous autoregressive (SAR)–type models.
A wide range of SAR-type models exists, including the spatial error model (SEM), the spatial autoregressive model (SAM), and the spatial Durbin model (SDM), the latter of which incorporates spatially lagged covariates (Anselin,, 1988; Cressie,, 1993). The selection model factorisation is a highly general decomposition and can be applied to any of these SAR-type models with only minor modifications. For illustration, consider the SDM:
$\textbf{y}=\rho\textbf{W}\textbf{y}+\textbf{X}\boldsymbol{\beta}_{1}+\textbf{X}_{s}\boldsymbol{\beta}_{2}+\textbf{e},$ (S40)
where $\boldsymbol{\beta}_{1}=(\beta_{10},\dots,\beta_{1r})^{\top}$ is the $(r+1)\times 1$ vector of fixed effects associated with the covariates in the design matrix X. The $n\times r$ matrix $\textbf{X}_{s}=\textbf{W}\textbf{X}^{o}$ represents the spatially lagged covariates, where $\textbf{X}^{o}$ contains the covariates excluding the intercept column. The vector $\boldsymbol{\beta}_{2}=(\beta_{21},\dots,\beta_{2r})^{\top}$ is the $r\times 1$ vector of fixed effects corresponding to the spatially lagged covariates, and $\textbf{e}\sim N(0,\sigma^{2}_{\textbf{e}}\textbf{I}_{n})$.
The parameter vector for the SDM with Gaussian errors is therefore $\boldsymbol{\phi}=(\boldsymbol{\beta}_{1}^{\top},\boldsymbol{\beta}_{2}^{\top},\rho,\sigma^{2}_{\textbf{e}})^{\top}$}. In the Gaussian case, $\boldsymbol{\xi}=\boldsymbol{\phi}$. If we instead assume Student’s-$t$ errors, $e_{i}\sim t_{\nu}(0,\sigma^{2}_{\textbf{e}})$, then the parameter vector becomes $\boldsymbol{\phi}=(\boldsymbol{\beta}_{1}^{\top},\boldsymbol{\beta}_{2}^{\top},\rho,\sigma^{2}_{\textbf{e}},\nu)^{\top}$}, and $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, which is consistent with the extension of the SEMs to Student’s-$t$ errors (SEM-t) described in Section 2.1 of the main paper.
Importantly, the same selection model factorisation used in Equation (13) of the main paper applies directly to both the Gaussian SDM and its Student’s-$t$ extension, with the only difference being the form of the parameter vector $\boldsymbol{\xi}$. Specifically, the joint density of the missingness indicators m and the response vector y is factorised as
$p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})=p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})\,p(\textbf{y}\mid\boldsymbol{\xi}).$ (S41)
Thus, the methodological framework developed in the main paper, namely, extending SEMs to non-Gaussian settings, jointly modelling the missingness indicators and the spatial process via the selection model factorisation, and performing inference using the HVB algorithm, is fully general and can be readily adapted to other SAR-type models with alternative spatial dependence structures.
S5 Additional derivations for the HVB algorithm
This section provides additional proofs and derivations for the proposed HVB algorithm used to estimate non-Gaussian SEMs with missing data, introduced in Section 4.2 of the main paper.
S5.1 Derivation of the reparameterisation gradient
@@
Since $\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{q_{\boldsymbol{\lambda}}}\left[\text{log}~p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda})$, as shown in Equation (19) of the main paper, the reparameterisation gradient of $\mathcal{L}$ is the same as that of $\mathcal{L}^{0}$ and given by